# Disentangle and Regularize: Sign Language Production with Articulator-Based Disentanglement and Channel-Aware Regularization

### v2

@@ 4.1 Datasets
We evaluate our model on two continuous sign language datasets. PHOENIX-2014T [3] contains 8,247 German Sign Language sentences aligned with gloss and spoken German. We use the 3D pose annotations from the CVPR 2025 SLRTP Challenge [31], where 2D keypoints extracted via MediaPipe Holistic [16] are uplifted to 3D using Ivashechkin et al. ’s method [13], yielding $N\times 178\times 3$ tensors. CSL-Daily [35] includes 20,654 Chinese Sign Language (CSL) sentences from 10 signers. We extract 3D poses using MediaPipe Holistic, standardizing each frame to 178 keypoints (upper body, hands, face). To ensure consistency across signers and datasets, poses are normalized by centering on the neck and scaling by shoulder width.
4.2 Training Procedure
Our training strategy follows a two-stage process to effectively map spoken language sentences to continuous sign pose sequences. In the first stage, we train a structurally disentangled pose autoencoder to learn compact multi-modal latent representations of sign language poses. The autoencoder is optimized to accurately reconstruct the original poses while ensuring meaningful decomposition of manual (hands) and non-manual (face, upper body) articulators. In the second stage, we train a non-autoregressive transformer to predict these latent pose representations conditioned on the input text. This stage utilizes sentence embeddings and is supervised with the ground truth encodings produced by the pretrained pose encoder.
4.2.1 Training the Pose Autoencoder
The training objective of our structurally disentangled autoencoder is to minimize the weighted reconstruction loss across four distinct anatomical regions: upper body, right hand, left hand, and face. The reconstruction loss for each region is computed independently using the L1 loss function, reflecting the absolute difference between the input keypoints and their reconstructions. The total loss is computed as the weighted sum of these individual components. The weights are set to 1.5 for both hands, reflecting their critical role in manual articulation, and 0.5 and 1.0 for the upper body and face, respectively. The reason we down-scale the weight of upper body joints is to prevent the loss function from being dominated by high-variance, large-scale motions, which can lead the model to overlook fine-grained but semantically important hand and facial movements. The total loss is provided in (1).
$\mathcal{L}_{recon}=\sum_{r\in\{\text{B},\,\text{RH},\,\text{LH},\,\text{F}\}}w_{r}\frac{1}{N_{r}}\sum_{i=1}^{N_{r}}\|\hat{\mathbf{x}}_{r}^{(i)}-\mathbf{x}_{r}^{(i)}\|_{1}\$ (1)
where; $\mathbf{x}_{r}^{(i)}$ and $\hat{\mathbf{x}}_{r}^{(i)}$ denote the ground truth and reconstructed keypoints for region $r$ at frame $i$, $N_{r}$ is the total number of frames, $w_{r}$ is the weight assigned to region $r$, set empirically as $w_{\text{B}}=0.5$, $w_{\text{RH}}=1.5$, $w_{\text{LH}}=1.5$, and $w_{\text{F}}=1.0$. The regions B, RH, RL, F represent body, right hand, left hand and face channels, respectively.
$\mathcal{L}_{\text{reg-enc}}=\lambda\sum_{j\in\{\text{enc}\}}\|\mathbf{W}_{j}\|_{1}$ (2)
To encourage sparsity in the latent representations and prevent overfitting, we apply an L1 regularization term to the encoder weights. The regularization coefficient ($\lambda$) is set empirically to $1\times 10^{-4}$ based on cross-validation. Applying L1 regularization solely to the encoder, while leaving the decoder unregularized, is a strategy adopted in sparse autoencoders [18], [27]. By not constraining the decoder, we retain its full capacity to reconstruct the fine-grained spatial dependencies inherent in sign articulation, particularly in manual articulators such as fingers. In contrast, imposing sparsity on the decoder could unnecessarily limit its expressiveness and compromise reconstruction fidelity.
The total autoencoder loss function combines reconstruction and regularization terms:
$\mathcal{L}_{\text{total-ae}}=\mathcal{L}_{\text{recon}}+\mathcal{L}_{\text{reg-enc}}$ (3)
Optimization is performed using the Adam optimizer, with a learning rate of $2\times 10^{-4}$ and beta parameters set to $(0.5,0.9)$, ensuring stable convergence during training.
4.2.2 Training Generator Model
The training objective of our transformer model is divided into two phase to effectively balance the learning of accurate pose encodings and structured latent space distribution. In the first-phase, the model is trained to predict ground truth (GT) pose encodings obtained from the pretrained pose autoencoder. In this phase, we minimize an L1 loss between the predicted encodings for the transformer model and the obtained GT encodings. We will refer to this model as DARSLP.
For a sequence of length $T$, the total pose encodings loss is defined as:
$\mathcal{L}_{\text{enc}}=\sum_{R\in\{\text{B},\text{RH},\text{LH},\text{F}\}}w_{R}\frac{1}{T}\sum_{t=1}^{T}\left\|\hat{z}_{t}^{(R)}-z_{t}^{(R)}\right\|_{1}$ (4)
where $\hat{z}_{t}^{(R)}$ and $z_{t}^{(R)}$ are the predicted and GT encodings at timestep $t$, and $w_{R}$ denotes region-specific weights. We empirically set $w_{\text{B}}=1$, $w_{\text{RH}}=7$, $w_{\text{LH}}=5$, $w_{\text{F}}=2$, emphasizing the importance of manual articulators.
Since hands carry the core lexical content in sign language, they are given the highest weights ($w_{\text{RH}}=7$, $w_{\text{LH}}=5$). The right hand, in particular, is typically the dominant hand in sign languages like DGS and CSL, and is primarily responsible for conveying meaning, while the left hand often plays a supportive or symmetrical role. The body provides broader spatial cues but is less semantically dense per frame, so it has a lower weight ($w_{\text{B}}=1$). The face, while critical for grammatical and emotional expressions, involves highly redundant motion across many keypoints, leading to a moderate weight ($w_{\text{F}}=2$). This design ensures the model focuses on the most informative regions for sign language generation.
In addition, the model predicts the relative sequence length, supervised by an L1 length loss:
$\mathcal{L}_{\text{len}}=\left\|\hat{r}-r\right\|_{1}$ (5)
where, $\hat{r}$ is the predicted length ratio, and $r$ is the ground truth. The overall objective in this phase becomes:
$\mathcal{L}_{\text{phase1}}=\mathcal{L}_{\text{enc}}+\mathcal{L}_{\text{len}}$ (6)
In the second phase of transformer training, we incorporate a KL divergence term to promote generalization and preserve the distributions of the structurally disentangled latent space. We refer to the resulting model as DARSLP-KL, which extends the first-phase model by adding regularization on the prior channel distributions. In this context, during training, we enforce the predicted latent channel distributions to align with precomputed priors derived from ground truth encodings for each articulator region. This loss term is provided in Eqn. (7).
$\mathcal{L}_{\text{KL}}=\sum_{R}\lambda_{R}\sum_{c\in R}D_{\text{KL}}\left(\mathcal{N}(\mu^{(c)},\sigma^{(c)})\,\|\,\mathcal{N}(\mu_{\text{prior}}^{(c)},\sigma_{\text{prior}}^{(c)})\right)$ (7)
Here, $\mu^{(c)}$ and $\sigma^{(c)}$ denote the batch-wise mean and variance of the predicted encoding channels, computed dynamically at each training iteration, while $\mu_{prior}^{(c)}$ and $\sigma_{prior}^{(c)}$ represent the prior statistics derived from the training data embeddings. Here, $R$ is the similar region set $\{B,RH,LH,F\}$ that we explained previously, and each channel is weighted according to the region it belongs to. For initial experiments we set the weights as $\lambda_{\text{B}}=1$, $\lambda_{\text{RH}}=5$, $\lambda_{\text{LH}}=3$, $\lambda_{\text{F}}=1$.
The total loss in this phase becomes as in Eqn. (8).
$\mathcal{L}_{\text{phase2}}=\mathcal{L}_{\text{enc}}+\mathcal{L}_{\text{len}}+\mathcal{L}_{\text{KL}}$ (8)
@@ 4.2.2 Training Generator Model
We employ the Adam optimizer with a learning rate of $2\times 10^{-4}$, weight decay of $1\times 10^{-4}$, and a ReduceLROnPlateau scheduler (factor 0.9, patience 40). Early stopping based on validation loss is applied to prevent overfitting. Our model is implemented using PyTorch Lightning [5], trained on a HPC with 4 NVIDIA A100-SXM4-40GB GPU.
4.3 Ablations
Ablation study on latent dimensionality. To assess the impact of latent dimensionality on sign pose generation, we conduct an ablation study by varying the allocation of dimensions across body, hands, and face components. These experiments test whether increased capacity improves downstream performance and validate our design choices. We compare three configurations: (1) 80 dimensions (64 for body and hands, 16 for face), (2) 96 dimensions (64 + 32), and (3) 160 dimensions (128 + 32). Within each, body and hand dimensions are distributed proportionally to joint counts to maintain anatomical consistency.
As shown in Table 1, the 80-dimensional setup achieves the best overall performance. Increasing face capacity to 32 dimensions shows no gain, while expanding body and hand dimensions to 128 degrades results. The ablation further validates our face encoding decision. We demonstrated that the 128×3 facial keypoints, due to their strong inter-keypoint dependency, can be effectively compressed to just 16 latent dimensions without compromising accuracy. This structural redundancy in facial movements, confirmed by our variance analysis enables robust and efficient encoding (see Section 5.1). Increasing the dimensionality beyond this point introduces unnecessary capacity without a gain, confirming the compact configuration as both sufficient and optimal.
Table 1: Performance comparison of the models with different latent dimensions on the PHOENIX14T DEV Set. All models are trained without KL divergence.
Models DTW-MJE↓ BLEU-1↑ BLEU-2↑ BLEU-3↑ BLEU-4↑ chrF↑ ROUGE↑
GT 0.000 30.92 21.34 16.01 12.74 34.71 30.19
128 (Body+Hands) + 32 (Face) 0.042 22.45 13.78 9.58 7.21 28.18 20.70
64 (Body+Hands) + 32 (Face) 0.044 20.92 12.57 8.79 6.72 27.78 19.91
64 (Body+Hands) + 16 (Face) 0.041 22.59 13.87 9.83 7.59 28.25 21.19
Ablation study on generator training. To determine the optimal setup for our transformer-based generator, we conducted ablation studies on training strategies and region-wise weighting. We initially used moderate region weights (1–7–5–2) for the L1 loss and (1–3–5–1) for the KL divergence, corresponding to the body, right hand, left hand, and face regions, respectively, to reflect articulator importance without imposing a strong bias. This served as a baseline for assessing the impact of KL regularization and its scheduling. As shown in Table 2, omitting KL leads to suboptimal performance (BLEU-4: 7.59), while applying it from the start causes instability (BLEU-4: 6.93). Our two-stage training strategy, introducing KL after initial training, yields the best results (BLEU-4: 8.70). This aligns with findings in VAE-based models, where premature KL application can collapse the latent space toward the prior [2], prompting staged or cyclical KL schedules [25, 6]. Though our model does not learn a variational posterior, our findings suggest that latent space regularization benefits from careful scheduling of KL divergence even in deterministic, structured representation settings.
After establishing the training pipeline with this balanced setup, we performed further ablations focusing on region-specific weighting. These experiments explore how emphasizing specific articulators (particularly the hands) in both L1 and KL terms influences the generation quality (Tables 3 and 4).
Table 2: Performance comparison of the models with different training strategies on the PHOENIX14T DEV Set. All models use 80-dimensional latent space.
Models DTW-MJE↓ BLEU-1↑ BLEU-2↑ BLEU-3↑ BLEU-4↑ chrF↑ ROUGE↑
GT 0.000 30.92 21.34 16.01 12.74 34.71 30.19
L1 w/o KL 0.041 22.59 13.87 9.83 7.59 28.25 21.19
L1 + w/ KL 0.042 20.51 12.46 8.88 6.93 27.00 18.95
Pretrained(w/L1)+Tuning w/ KL 0.039 25.91 16.34 11.43 8.70 29.92 23.67
Table 3: Performance comparison of models with different region weight configurations on the PHOENIX14T DEV set. All models use 80-dimensional latent space, and are trained without KL regularization.
Weighting Strategy (Body–RH–LH–Face) DTW-MJE↓ BLEU-1↑ BLEU-2↑ BLEU-3↑ BLEU-4↑ chrF↑ ROUGE↑
Uniform (1–1–1–1) 0.039 22.14 13.14 8.95 6.72 27.89 19.91
Moderate Hand Emphasis (1–7–5–2) 0.041 22.59 13.87 9.83 7.59 28.25 21.19
Heavy Hand Emphasis (1–14–10–2) 0.043 24.78 15.84 11.26 8.66 29.36 22.98
Table 4: Ablation results with different KL divergence weights with fixed L1 weighting (1–14–10–2) on the PHOENIX14T DEV set. All models use 80-dimensional latent space, and are trained with staged KL regularization.
KL Weights (Body–RH–LH–Face) DTW-MJE↓ BLEU-1↑ BLEU-2↑ BLEU-3↑ BLEU-4↑ chrF↑ ROUGE↑
Uniform (1–1–1–1) 0.040 26.69 17.49 12.79 9.99 30.63 24.92
Moderate Hand Emphasis (1–3–5–1) 0.040 27.35 17.51 12.52 9.66 30.96 25.28
Heavy Hand Emphasis (1–10–6–1) 0.039 27.11 18.00 13.09 10.27 31.38 25.70
Based on the ablation results, region-wise L1 loss weights are set to (1–14–10–2), and KL divergence is introduced in the second phase with weights (1–10–6–1).
5 Discussions
5.1 Analysis of the Learned Latent Spaces
(a) Top 8 right hand channels with highest entropy. Average entropy across all channels: 2.0101.
(b) Top 8 left hand channels with highest entropy. Average entropy across all channels: 2.2241.
(c) Top 8 face channels with highest entropy. Average entropy across all channels: 0.1377.
(d) Top 8 body channels with highest entropy. Average entropy across all channels: 1.1075.
Figure 3: Histograms of top 8 latent channels with the highest entropy for each region. Each subplot shows the distribution of a channel along with its entropy (E), interquartile range (IQR), and standard deviation (SD). (Zoom in for better visibility.)
@@ 5.1 Analysis of the Learned Latent Spaces
The statistical analysis of the learned latent representations reveals meaningful distinctions in the encoding behavior across body regions. In Figure 3, we present example histograms and corresponding entropy values for selected channel distributions from each structurally disentangled latent subspace (face, body, right hand, and left hand) learned by the autoencoder.

### v3

@@ 4.1 Implementation Details
Datasets. We evaluate our model on two continuous sign language datasets. PHOENIX-2014T [3] contains 8,247 German Sign Language sentences aligned with gloss and spoken German. We use the 3D pose annotations from the CVPR 2025 SLRTP Challenge [31], where 2D keypoints extracted via MediaPipe Holistic [16] are uplifted to 3D using Ivashechkin et al. ’s method [13], yielding $N\times 178\times 3$ tensors. CSL-Daily [36] includes 20,654 Chinese Sign Language (CSL) sentences from 10 signers. We extract 3D poses using MediaPipe Holistic, standardizing each frame to 178 keypoints (upper body, hands, face). To ensure consistency across signers and datasets, poses are normalized by centering on the neck and scaling by shoulder width.
Autoencoder training settings. We assign region-specific weights to the reconstruction loss: $w_{\text{RH}}=w_{\text{LH}}=1.5$, $w_{\text{F}}=1.0$, $w_{\text{B}}=0.5$. These weights prevent high-variance upper body motions from overshadowing fine-grained hand and facial cues. L1 regularization with $\lambda=1\times 10^{-4}$ promotes encoder sparsity. Optimization is performed using the Adam optimizer, with a learning rate of $2\times 10^{-4}$ and beta parameters set to $(0.5,0.9)$. The pose autoencoder is trained for 270 epochs on both PHOENIX14T and CSL-Daily datasets.
Transformer training settings. For L1 loss, we use; $w_{\text{RH}}=14$, $w_{\text{LH}}=10$, $w_{\text{F}}=2$, decided based on ablation studies (Section 4.3) comparing different weighting schemes. Hands carry the core lexical content in sign language, with the right hand as dominant in DGS and CSL, conveying most meaning, and the left hand serving a supportive role. The body offers coarse spatial context, while the face encodes grammatical and emotional cues but involves redundant motion across many keypoints. We apply similar region-based weighting in both the autoencoder and transformer training settings to reflect these functional roles. Regions are weighted accordingly, guided by ablation results.
@@
We employ the Adam optimizer with a learning rate of $2\times 10^{-4}$, weight decay of $1\times 10^{-4}$, and a ReduceLROnPlateau scheduler (factor 0.9, patience 40) in both phases of the training. Early stopping based on validation loss is applied to prevent overfitting. Training is performed on a 4×NVIDIA A100-SXM4-40GB setup using PyTorch Lightning [5].
4.2 Ablation Study on Latent Dimensionality
We first analyze how different allocations of latent dimensions across body, hands, and face affect performance. Table 1 compares multiple configurations, keeping the body+hands dimension fixed or varied and adjusting the face block. Results are reported on the PHOENIX14T development set for all ablations.
Latent Dim Config. B1↑ B2↑ B3↑ B4↑ chrF↑ ROUGE↑
32 hands&body + 8 face 20.79 11.81 8.34 6.43 27.01 18.44
64 hands&body + 8 face 22.54 13.84 9.75 7.44 28.58 21.18
128 hands&body + 8 face 20.44 11.99 8.45 6.54 27.43 18.67
64 hands&body + 16 face 24.07 14.25 9.73 7.25 28.51 21.30
64 hands&body + 32 face 23.04 14.00 9.88 7.54 28.67 21.42
64 hands&body + 64 face 20.75 12.39 8.72 6.67 27.71 19.07
80-dim entangled (no split) 19.61 11.34 7.97 6.13 27.21 18.10
Table 1: Effect of different latent dimension configurations.
The ablation results in Table 1 show that the best overall performance is achieved with 64 latent dimensions allocated to body and hands, distributed proportionally to joint counts (8 for body, 28 for each hand) to preserve anatomical consistency. For the face, despite having a much larger number of keypoints (128), we observe that 8, 16, and 32 latent dimensions all produce very similar results. This indicates strong interdependence among facial keypoints, allowing the face representation to be effectively compressed into a compact latent block without loss of information. Furthermore, all disentangled models outperform the entangled configuration with an 80-dimensional flat latent vector in BLEU-4 and ROUGE metrics, highlighting the benefit of factorized and structured representations.
4.3 Ablation Study on Generator Training
We next analyze the impact of region-specific L1 loss weighting on generator training. In this context, the three numbers in ($w_{R\text{H}}-w_{L\text{H}}-w_{F}$) indicate the weighting coefficients assigned to the right hand, left hand, and face regions, respectively. As shown in Table 2, moderate weighting of the hands, such as (5-5-1), consistently improves performance over the unweighted baseline (1-1-1), demonstrating the benefit of emphasizing manual components. Further gains are achieved by prioritizing the dominant right hand, with the best results observed at the (14-10-2) configuration for the 64H&B + 16F setup. Increasing the weights beyond this point causes a clear performance drop.
RH-LH-F B1↑ B2↑ B3↑ B4↑ chrF↑ ROUGE↑
64H&B + 32F
(1-1-1) 23.04 14.00 9.88 7.54 28.67 21.42
(5-5-1) 23.64 14.90 10.50 8.13 29.52 22.30
(7-5-1) 20.76 12.25 8.44 6.46 27.51 19.36
(14-10-2) 20.55 12.58 8.86 6.79 27.19 18.93
(21-15-2) 19.92 12.20 8.50 6.40 26.72 18.12
64H&B + 16F
(1-1-1) 24.07 14.25 9.73 7.25 28.51 21.30
(7-5-1) 23.96 15.16 10.68 8.06 29.23 22.48
(14-10-1) 24.18 15.47 11.00 8.50 29.52 22.81
(14-10-2) 24.97 15.91 11.35 8.72 29.67 23.25
(14-10-4) 22.85 13.82 9.49 7.14 27.91 21.21
(21-15-2) 24.51 15.18 10.59 7.88 29.27 22.30
64H&B + 8F
(1-1-1) 22.54 13.84 9.75 7.44 28.58 21.18
(5-5-1) 23.03 14.46 10.38 8.03 29.08 21.62
(7-5-1) 22.18 13.73 9.53 7.19 27.90 20.78
(14-10-2) 21.59 13.47 9.44 7.16 27.47 19.27
(21-15-2) 20.98 12.88 8.90 6.80 26.84 18.70
Table 2: Effect of varying L1 loss weights for right hand (RH), left hand (LH), and face (F) components while keeping body+hands fixed at 64 dimensions. Bold entries indicate optimal configurations before degradation sets in.
We further investigate the impact of KL regularization and its scheduling. As shown in Table 3, omitting KL results in suboptimal performance, while applying it from the start causes instability. Our two-phase strategy, where the model is first trained with only L1 loss and then fine-tuned with KL, achieves the best results across all metrics. This supports prior findings in VAE literature [2, 25, 6] that KL regularization must be carefully scheduled. While our model does not follow a variational formulation, the findings emphasize that applying regularization in a staged manner can still play a critical role in enforcing structure within the latent space.
Model B1↑ B2↑ B3↑ B4↑ chrF↑ ROUGE↑
GT (dev) 30.92 21.34 16.01 12.74 34.71 30.19
L1 w/o KL 24.97 15.91 11.35 8.72 29.67 23.25
L1 + w/ KL 20.78 12.64 9.01 6.98 27.92 19.66
Two-Phase (Ours) 28.41 19.31 14.45 11.40 31.85 26.93
Table 3: Effect of KL divergence on training strategies using 80-dim latent space and (14-10-2) L1 weighting.
4.4 Latent Space Analysis
To isolate and interpret the contribution of each articulator within the latent spaces of both entangled and disentangled autoencoders, we adopt a targeted visualization strategy on the PHOENIX14T dataset. For each input sample, only the keypoints of a selected region (e.g., right hand) are retained, while all other keypoints are replaced with a canonical reference pose that is computed as the mean pose over the training set. This masking ensures that variation in the latent space is driven solely by the selected region, while the remaining parts of the body remain fixed, minimizing confounding effects and providing a stable reference frame. The masked sequences are then passed through the model to obtain 80-dimensional latent representations, which are standardized using training statistics for visualization.
Principal Component Analysis (PCA) [33] is first applied to the latent representations to suppress noise and retain the most informative variance components. This is followed by Uniform Manifold Approximation and Projection (UMAP) [18], which maps the PCA-reduced features to a 2D space for clearer visualization. Both PCA and UMAP are fitted on the training set and consistently applied to the development set to ensure comparability. This unified analysis pipeline allows fair comparison between entangled and disentangled autoencoders. In particular, masking is crucial for the entangled AE, as it does not maintain region-specific latent channels and would otherwise produce entangled activations influenced by irrelevant articulators.
In the entangled AE visualization, the body and face are represented as compact, near-unimodal clusters with substantial overlap between different articulators (Fig. 3(a)). In contrast, the disentangled AE preserves the intrinsic variance of each region, resulting in more structured and well-separated latent distributions 3(b)). The hands and face, in particular, display rich and complex multi-modal patterns that reflect the fine-grained dynamics of sign language movements. This separation reveals underlying structures in the data that are suppressed in the entangled representation, indicating that the disentangled AE is more effective at capturing region-specific motion patterns critical for sign language production.
(a) Entangled AE
(b) Disentangled AE
Figure 3: Masked UMAP comparison of entangled vs. disentangled AE. Colors represent different regions; gray shows unmasked latents.
In addition to analyzing the latent space of autoencoders, we further examine how the latent representations produced by our Transformer-based generator compare to those of the AE and how KL regularization influences these representations. To this end, we perform a separate PCA analysis for each articulator. Latent channels corresponding to the body, right hand, left hand, and face are standardized using region-specific training statistics, and distinct PCA models are fitted on each region’s training data to obtain independent 2D projections.
As shown in Fig. 4(a), the AE produces compact, stable distributions while transformer without KL regularization produces more scattered and noisy patterns that deviate from the AE (Fig. 4(b)). Introducing KL regularization guides the Transformer’s distributions toward the AE prior, reducing noise and enhancing compactness, particularly in the face and hand regions 4(c), indicating that KL improves the structural organization of the latent space.
(a) GT regional PCA
(b) Transformer w/o KL
(c) Transformer w/ KL
(d) GT - Transformer w/o KL
(e) GT - Transformer w/ KL
Figure 4: (a–c) Regional PCA projections. (d-e) Density-difference maps (GT – Transformer): red indicates areas with higher density than the AE, blue indicates lower.
To quantify the mismatch between the AE and the Transformer, we compute smooth 2D density estimates within each PCA plane and visualize the difference (GT minus Transformer) as heatmaps. Fig. 4(c–d) illustrates the discrepancies between the two projected distributions. Without KL regularization, deviations are diffuse and scattered; particularly in the face and hand regions. When KL is applied, these differences become more centered and structured, suggesting that KL promotes a more coherent alignment of the Transformer’s latent space with the AE prior.
4.4.1 Statistical Analysis of the Latent Spaces
@@
Further, statistical analysis of the learned latent representations reveals meaningful distinctions in the encoding behavior across regions for both dataset. In Figure 5, we present example histograms and corresponding entropy values for selected channel distributions from each structurally disentangled latent subspace (face, body, right hand, and left hand) learned by the autoencoder.
(a) Right hand channels (avg. entropy: 2.01).
(b) Left hand channels (avg. entropy: 2.22).
(c) Face channels (avg. entropy: 0.14).
(d) Body channels (avg. entropy: 1.11).
Figure 5: Histograms of top 8 latent channels with the highest entropy for each region of PHOENIX14T Dataset. Each subplot shows the distribution of a channel along with its entropy (E), interquartile range (IQR), and standard deviation (SD). (Zoom in for better visibility.)